OpenStudy (anonymous):

Find x3 - 2y2 - 3x3 + z4 if x = 3, y = 5, and z = -3.

OpenStudy (baby_bear69):

\(\Huge\bf\color{#FF0000}W\color{#FF4900}e\color{#FF9200}l\color{#FFDB00}c\color{#FFff00}o\color{#B6ff00}m\color{#6Dff00}e\color{#24ff00}~\color{#00ff00}t\color{#00ff49}o\color{#00ff92}~\color{#00ffDB}O\color{#00ffff}p\color{#00DBff}e\color{#0092ff}n\color{#0049ff}S\color{#0000ff}t\color{#2400ff}u\color{#6D00ff}d\color{#B600ff}y\color{#FF00ff}!\)

OpenStudy (texaschic101):

x^3 - 2y^2 - 3x^3 + z^4 ---- x = 3, y = 5, z = -3 now we sub 3^3 - 2(5^2) - 3(3^3) + (-3^4) can you do the math now ?

OpenStudy (anonymous):

-185

OpenStudy (texaschic101):

27 - 2(25) - 3(27) + 81 27 - 50 - 81 - 81 -212 + 27 -185 thats what I got

OpenStudy (texaschic101):

oops..made a mistake on first line...supposed to be -81

OpenStudy (anonymous):

yep

OpenStudy (texaschic101):

same answer....-185

OpenStudy (anonymous):

so it is -185

OpenStudy (texaschic101):

yes

OpenStudy (anonymous):

i got it wrong

OpenStudy (texaschic101):

hold on...let me check again 3^3 - 2(5^2) - 3(27) + 81 ---- it is + 81 because -3^4 = -3*-3*-3*-3 = 81 27 - 50 - 81 + 81 - 23 (the 81's cancel each other out)

OpenStudy (texaschic101):

I am so sorry about that

OpenStudy (anonymous):

can i ask you another question

OpenStudy (anonymous):

??

OpenStudy (anonymous):

Find -3a2 - b3 + 3c2 - 2b3 if a = 2, b = -1, and c = 3.